OpenStudy (anonymous):

I don't know how to solve this bc 1 isn't a fraction. Which of these numbers is between 1 and 3/5? A. 8/5 B. 5/9 C. 5/8 D. 3/8

OpenStudy (anonymous):

One can be equal to 1.0 and 3/5 can be equal to .6. Convert all the choices to decimals to get your answer. A. 8/5=1.6 B. 5/9=about .55 C.5/8=.625 D. 3/8=.375 .625 is the only decimal in between 1 and .6, therefore C is your answer.

OpenStudy (anonymous):

Thanks. What if 3/5 became 3/4? @Sionainn

OpenStudy (anonymous):

3/4=.75. None of the answers are in between .75 and 1, so there would be no correct answer unless the options changed. And I'm glad I could help :)

OpenStudy (anonymous):

Oh they do, A. 7/4 B. 7/8 C. 3/5 D. 5/8 @Sionainn

OpenStudy (anonymous):

@zpupster

OpenStudy (texaschic101):

turn all the fractions to decimals 7/4 = 1.75 7/8 = 0.875 3/5 = 0.6 5/8 = 0.625 so what do you think the answer is ?

OpenStudy (anonymous):

1 is 1.0 and 3/4 is 0.75. So B? @texaschic101

OpenStudy (texaschic101):

you got it...it's B

OpenStudy (anonymous):

thank you so much!!!! @texaschic101 @Sionainn

OpenStudy (anonymous):

Glad I could help :)

OpenStudy (texaschic101):

there is a couple of ways to do problems such as these...either you can turn them all into fractions with the same denominators, or you can turn them to decimals. Either way, you should arrive at the same answer. Glad I could help :)